If the mitochondrial matrix has a pH of 8.4, what is the number of protons inside a respiring mitochondrion from liver? the matrix compartment of the mitochondrion is a sphere with a diameter of 1.6 micro meter.
Answer #1

pH=-log[H+]

pH=8

8=-log[H+]

[H+]=antilog (-8)

[H+]=10^-8M

volume of the sphere=4/3r^3

matrix compartment of the mitochondrion is a sphere with a diameter of 1.6 micro meter.

r=diameter/2=1.6/2=0.8 um

volume of the mitochondrial matrix=4/3 (0.8)^3

= 2.144 um^3

1 um=10^-6 m

so 1 um^3= (1 10^-6)^3=10^-18 m^3

so 2.144 um^3= 2.144 10^-18 m^3

1m^3=1000L

so 2.144 10^-18 m^3=2.144 10^-181000=2.144 10^-15 L

concentration of H+= 10^-8 M

molarity= number of moles/volume of the solution in Litres

number of moles= molarity volume of the solution in litres

= 10^-8 2.144 10^-15

= 2.144 10^-23

1 mole contains 6.02210^23 molecules

so number of protons in the mitochondrial matrix= number of moles of H+ 6.02210^23

= 2.144 10^-23 6.02210^23

= 14.7

number of protons in the mitochondrial matrix= 14.7